OpenStudy (anonymous):

The function f(t) = 65 sin (pi over 5t) + 35 models the temperature of a periodic chemical reaction where t represents time in hours. What are the maximum and minimum temperatures of the reaction, and how long does the entire cycle take? Maximum: 65°; minimum: 35°; period: pi over 5 hours Maximum: 100°; minimum: −30°; period: pi over 5 hours Maximum: 65°; minimum: 35°; period: 10 hours Maximum: 100°; minimum: −30°; period: 10 hours

OpenStudy (anonymous):

This i dont have a clue on. xD

OpenStudy (anonymous):

lol which one you want to do first, max, min or period? both are very easy

OpenStudy (anonymous):

Period ;o

OpenStudy (anonymous):

Wait, dont they already give us the period? pi/5?

OpenStudy (anonymous):

on no!!

OpenStudy (anonymous):

i know what you are thinking, that the period of \[\sin(bx)\] is \(b\) but it is not the period of \[\sin(bx)\] is \[\frac{2\pi}{b}\]

OpenStudy (anonymous):

OH. Then the Period is 10 ;o

OpenStudy (anonymous):

you are quicker on the algebra than i am yes, 10 easy right?

OpenStudy (anonymous):

now we need the range

OpenStudy (anonymous):

So, Its either C or D c:

OpenStudy (anonymous):

that is also easy do you know what the biggest sine can be? just asking, if not let me know and i will tell you

OpenStudy (anonymous):

65?

OpenStudy (anonymous):

lets go slow

OpenStudy (anonymous):

the largest sine can be is 1 so we can replace \[\sin(\frac{\pi}{5}t)\] by \(1\) and compute \[65\times 1+35\] a pretty fast computation gives \(100\) as the max

OpenStudy (anonymous):

OH. So its D :3

OpenStudy (anonymous):

the smallest it can be is \(-1\) and \[65\times (-1)+35=-30\]

OpenStudy (anonymous):

right, D

OpenStudy (anonymous):

Youre too great c: can i have your input on an answer i put?